OpenStudy (anonymous):

Solve the triangle. A = 52°, b = 14, c = 6

OpenStudy (zale101):



OpenStudy (zale101):

What law are you going to use?

OpenStudy (zale101):

Law of sine will pretty much get you the answer. Any ideas of how to use the law of sine?

OpenStudy (anonymous):

sin?

OpenStudy (zale101):

Yes

OpenStudy (anonymous):

will A= 14.9 C=24.2 B=103.8 ?

OpenStudy (zale101):

The Law of Sine or the Sine rule says: \(\Large \frac{Sin(A)}{a}=\frac{Sine(B)}{b}=\frac{Sin(C)}{c}\)

OpenStudy (zale101):

How did you get that answer @Abbs__ ?

OpenStudy (anonymous):

I just plugged it in

OpenStudy (zale101):

Can you show me your work?

OpenStudy (anonymous):

idk how to do that, is it wrong ?

OpenStudy (zale101):

I'm not going to give out the answer unless you participate in your question and start showing some work.

OpenStudy (anonymous):

okay I obviously did the work, how do I take a picture of it and send it.

jimthompson5910 (jim_thompson5910):

@Zale101 the law of sines won't work here yet. We need to know either B or C. To start off, use the law of cosines to solve for 'a'. After that, the law of sines can be used.

OpenStudy (zale101):

Thanks @jim_thompson5910 !!!

jimthompson5910 (jim_thompson5910):

sure thing

OpenStudy (anonymous):

I literally did that tho

OpenStudy (zale101):

@Abbs__ Jim is right, you start off by using the law of cosines, to know B or C because apparently you can't go straight ahead and do the law of sines because it wont work. Your question only shows A, then b and c. We can't have Sin(A)/a because we only know A and a is not giving, same goes for B and C. There's missing parts that's why we can't use the laws of sine. Thanks again, jim Step 1: Laws of Cosine Step 2: Laws of sine

OpenStudy (zale101):

@Abbs__ can you then show us what you did?

OpenStudy (anonymous):

I asked how to do that twice

jimthompson5910 (jim_thompson5910):

Law of Cosines a^2 = b^2 + c^2 - 2*b*c*cos(A) a^2 = 14^2 + 6^2 - 2*14*6*cos(52) ... make the proper substitutions now isolate 'a'